OpenStudy (anonymous):

A high school student was able to save $8000 by working a part-time job every summer. He invested half the money in a money market account and half the money in a stock that paid three times as much interest as the money market account. After a year he earned $560 in interest. What were the interest rates of the money market account and the stock?

OpenStudy (anonymous):

lets call the interest rate of the money market \(x\) then the interest rate of the stock is \(3x\) you have \(4000\times x+4000\times 3x=560\) and you can solve for \(x\)

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

first step would be to combine like terms on the left and get \(1600x=560\) then divide to get the rate of interest

OpenStudy (anonymous):

oops i am off by a zero, should be \[16000x=560\] \[x=560\div16000=.035=3.5\%\]

OpenStudy (anonymous):

in addition: The Interest rate in the Money Market Account was _____ % The Interest rate in the Stock was_____________ %

OpenStudy (anonymous):

we solved for "money market" so the answer is \(3.5\%\)

OpenStudy (anonymous):

stock market is 3 times that amount, or \(10.5\%\)

OpenStudy (anonymous):

ist, 3(the anser) so i could get the interrest rate for stck?

OpenStudy (anonymous):

@satellite73 thank u. u're a real genius. ur so kind. thanks. =)

OpenStudy (anonymous):

yw (blush)